OpenStudy (anonymous):

Write an equation for the line that is parallel to the given line and passes through the given point. y = x + 8; (2, 16)

OpenStudy (anonymous):

y = x – 6 is this it?

OpenStudy (souvik):

any line that is parallel to y=x+8 has an equation y=x+ k where k is some constant and it passes through (2,16)..put these number in place of x and y

OpenStudy (anonymous):

16=2+K or 16=10

OpenStudy (souvik):

16=2+k or k=14

OpenStudy (souvik):

so the equation is y=x+14

OpenStudy (anonymous):

not one of my multible choices

OpenStudy (anonymous):

Since the line is parallel to y = x + 8, equating it with y=mx+c we find slope m=1(since the coefficient of x is 1) Now line passes through point (2, 16) . Thus \[x_{1} =2, y _{1}=16\] Eq of line in slope point form is given as \[y-y _{1}=m(x-x _{1})\] I.e. y- 16=1(x-2) I.e. y- 16=x-2 I.e. y=x-2 + 16 I.e. y=x+14 Which is the required eq of the line

OpenStudy (anonymous):

A. y = x B. y = x + 6 C. y = x – 6 D. y = x + 6 this is my choices

OpenStudy (souvik):

is the point is (2,16)?

OpenStudy (souvik):

B and D of your option is same..

OpenStudy (anonymous):

that is what it is on my home work

OpenStudy (anonymous):

is the point is (2,16)? yes that is what it says

OpenStudy (souvik):

hmm...

OpenStudy (anonymous):

thanks for your help

OpenStudy (souvik):

welcome...